Calculate the number of atoms present per cubic centimeters of zirconia (ZrO2), which has a density of
6.2 g/cm3
1
Expert's answer
2020-07-20T06:24:09-0400

Number of atoms can be calculate by the following formula:

N = n × NA

where N - number of atoms, n - number of moles, NA - Avogadro's number (6.022 × 1023).

Next, number of moles equals:

n = m / Mr

where m - mass, Mr - molecular weight (Mr(ZrO2) = 123 g/mol)

Finally, mass can be found as:

m = d × V

where d - density, V - volume

As a result, the final formula is:

N = n × NA = (m / Mr) × NA = (d × V / Mr) × NA = (6.2 g/cm3 × 1 cm3 / 123 g/mol) × 6.022 × 1023 = 3.035 × 1022


Answer: 3.035 × 1022